\begin{thebibliography}{14}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Kobayashi and Nomizu(1963)]{KobayashiNomizu1963fdg}
Shoshichi Kobayashi and Katsumi Nomizu.
\newblock \emph{Foundations of Differential Geometry}, volume~I.
\newblock New York, London: Interscience Publishers, John Wiley and Sons, 1963.

\bibitem[Hermann(1975)]{hermann1975g}
Robert Hermann.
\newblock \emph{Gauge fields and {C}artan-{E}hresmann Connections, Part A}.
\newblock Brookline, MA: Math Sci Press, 1975.

\bibitem[Nakahara(1990)]{nakahara1990gtp}
Mikio Nakahara.
\newblock \emph{Geometry, Topology, and Physics}.
\newblock Bristol and Philadelphia: Institute of Physics Publishing, 1990.

\bibitem[Carath\'eodory(1909)]{caratheodory1909}
Constantin Carath\'eodory.
\newblock Untersuchungen \"uber die {G}rundlagen der {T}hermodynamik.
\newblock \emph{Mathematische Annalen}, 67:\penalty0 355--386, 1909.

\bibitem[Bernstein(1960)]{bernstein1960c}
B.~Bernstein.
\newblock Proof of {C}arath\'eodory's local theorem and its global application
  to thermostatics.
\newblock \emph{Journal of Mathematical Physics}, 1\penalty0 (3):\penalty0
  222--224, 1960.

\bibitem[Boyling(1968)]{boyling1968c}
J.B. Boyling.
\newblock Carath{\'e}odory's principle and the existence of global integrating
  factors.
\newblock \emph{Communications in Mathematical Physics}, 10\penalty0
  (1):\penalty0 52--68, 1968.

\bibitem[Bryant et~al.(1991)Bryant, Chern, Gardner, Goldschmidt, and
  Griffiths]{BryantEtAl1991}
Robert~L Bryant, S.S. Chern, R.B. Gardner, H.L. Goldschmidt, and P.A.
  Griffiths.
\newblock \emph{Exterior Differential Systems}.
\newblock New York: Springer-Verlag New York Inc., 1991.

\bibitem[Jauch and B\'aron(1972)]{baronjauch1972a}
Josef~M. Jauch and J.G. B\'aron.
\newblock Entropy, information and {S}zilard's paradox.
\newblock \emph{Helvetica physica acta}, 45\penalty0 (2):\penalty0 220--232,
  1972.

\bibitem[Jauch(1972)]{jauch1972th}
Josef~M. Jauch.
\newblock On a new foundation of equilibrium thermodynamics.
\newblock \emph{Foundations of Physics}, 2\penalty0 (4):\penalty0 327--332,
  1972.

\bibitem[Jauch(1975)]{jauch1975at}
Josef~M. Jauch.
\newblock {Analytical Thermodynamics, Part 1: Thermostatics---General Theory}.
\newblock \emph{Foundations of Physics}, 5\penalty0 (1):\penalty0 111--132,
  1975.

\bibitem[Bravetti et~al.(2015)Bravetti, Lopez-Monsalvo, and
  Nettel]{bravetti-etal2015g}
Alessandro Bravetti, Cesar~S Lopez-Monsalvo, and Francisco Nettel.
\newblock Conformal gauge transformations in thermodynamics.
\newblock \emph{Entropy}, 17\penalty0 (9):\penalty0 6150--6168, 2015.
\newblock Open access, \url{https://dx.doi.org/10.3390/e17096150}.

\bibitem[Ferrari et~al.(2025)Ferrari, Rudnicki, and
  C{\'e}leri]{ferrarietal2025pra}
Gabriel~Fernandez Ferrari, {\L}ukasz Rudnicki, and Lucas~C C{\'e}leri.
\newblock Quantum thermodynamics as a gauge theory.
\newblock \emph{Physical Review A}, 111\penalty0 (5):\penalty0 052209, 2025.
\newblock \url{https://arxiv.org/abs/2409.07676}.

\bibitem[Frankel(2012)]{frankel2012gop}
Theodore Frankel.
\newblock \emph{The Geometry of Physics: An Introduction}.
\newblock Cambridge: Cambridge University Press, 3rd edition, 2012.

\bibitem[Shapere and Wilczek(1989)]{ShapereWilczek1989}
Alfred Shapere and Frank Wilczek.
\newblock \emph{Geometric Phases In Physics}.
\newblock Singapore, New Jersey, London, Hong Kong: World Scientific Publishing
  Co. Pte. Ltd., 1989.

\end{thebibliography}

\documentclass[twocolumn,a4paper,showkeys,amsmath,amssymb,prl]{revtex4-2}
\usepackage[dvipsnames]{xcolor} 
\usepackage{amsthm}
\usepackage{enumitem}
\usepackage{dsfont}
\usepackage{hyperref}
\hypersetup{
	colorlinks=true,         
	linkcolor=MidnightBlue,          
	citecolor=MidnightBlue,
	urlcolor=MidnightBlue            
      }
\usepackage{orcidlink}

\newtheorem{lemm}{Lemma}
\newtheorem{thm}{Theorem}
\newtheorem{coro}{Corollary}

\newcommand{\RR}{\mathds{R}} 

\DeclareMathOperator{\Aut}{Aut}
\usepackage{mathtools}
\def\dQ{{\mkern3mu\mathchar'26\mkern-12mu d}Q} 

\begin{document}

\title{Heat as a gauge connection}
\author{Bryan W. Roberts\,\orcidlink{0000-0003-0548-1280}
} 
 \email{b.w.roberts@lse.ac.uk}
 \affiliation{
Centre for Philosophy of Natural and Social Sciences, London School of Economics \& Political Science, Houghton Street, London WC2A 2AE, United Kingdom
} 

\date{\today}

\begin{abstract} 
We show that heat defines a gauge connection on a line bundle over work configurations. Vanishing curvature is equivalent to the local existence of entropy and temperature functions such that heat can be expressed as $TdS$. A conjecture of Jauch, that entropy and temperature arise from a conservation law, is shown to follow as a special case. Global equilibrium may nevertheless fail in the presence of a thermal analogue of geometric phase.
\end{abstract}
\keywords{Thermodynamics, gauge, holonomy, geometric phase, integrability, entropy, temperature}

\maketitle

\section{Introduction}\label{sec:intro}

It is well known that a gauge connection has vanishing curvature if and only if it is integrable \cite{KobayashiNomizu1963fdg,hermann1975g,nakahara1990gtp}, and that heat is integrable if and only if it can be locally expressed as $TdS$ \cite{caratheodory1909,bernstein1960c,boyling1968c,BryantEtAl1991}. In this paper we combine these two facts: viewing heat as a gauge connection, we show that vanishing thermal curvature is equivalent to the local existence of entropy and temperature functions. Our main result can be summarised as the statement: \emph{heat defines a connection on a line bundle over work configurations, which admits a local integrating factor whenever its curvature vanishes.} The proof of a conjecture of Jauch \cite{baronjauch1972a,jauch1972th,jauch1975at} regarding the existence of entropy and temperature follows as a special case. Moreover, the failure of global equilibrium in systems like the Wankel engine is shown to provide a formal analogue of geometric phase in gauge physics, such as the Berry phase. This provides new support for recent proposals to view thermodynamics as a special case of gauge physics \cite{bravetti-etal2015g,ferrarietal2025pra}. 

\section{The work bundle and gauge}

A physical system in thermodynamics is a smooth embedding $\varphi:L\to M$ of an $n$-dimensional manifold $L$ into a $(2n+1)$-dimensional manifold $M$. For example, the ideal gas with $n=2$ and global coordinates $(U,P,V,S,T)$ is constrained to a two-dimensional surface by the first law $TdS = dU - PdV$ together with the equations of state $PV = (nr)T$, and $U = f(T)$. We will show that such a surface $L$ can be naturally viewed as a line bundle with gauge symmetries.

A quantity of energy that can be extracted operationally using given devices, such as pistons or magnets, will be called `work'. In thermodynamics, work energy does not generally exhaust the total energy in a system, and so we refer to the remaining inaccessible energy as `heat'. Thus, let $N$ be the smooth manifold of dimension $n-1$ representing work configurations, and let total energy be a one-dimensional fibre over each such configuration. That is, let
\begin{equation}\label{eq:line-bundle}
  \pi:L\to N
\end{equation}
be a smooth submersion called the \emph{work line bundle}, where $L$ has dimension $n$.

Elegant expressions of energy, work, and heat are now available. First, let \emph{energy} be a vertical coordinate function $U:L\to\RR$, in that $dU(X)\neq0$ for all $X\in\ker\pi_*$. At this point, one could introduce local coordinates $(U, V_1,\dots, V_{n-1})$ for $L$ and define work along a curve in $L$ to be the one-form,
 \begin{equation}\label{eq:work}
   \omega := -\sum_{i=1}^{n-1}P_idV_i
 \end{equation}
where $P_i:L\to\RR$ is a smooth function for each $i=1,\dots,n-1$. However, our discussion will only need to make use of a weaker, coordinate-invariant property: let \emph{work} be any one-form that annihilates vertical vectors,
\begin{equation}\label{eq:work-kernel}
  \ker\pi_* \subseteq \ker\omega.
\end{equation}
This is true of $\omega$ in Equation \eqref{eq:work} because the vertical vector field $\partial/\partial U$ satisfies $dV_i(\partial/\partial U) = \partial V_i/\partial U = 0$ for each $i=1,\dots,n-1$. But, \eqref{eq:work-kernel} is a more direct way to capture the idea that the vertical directions are `invisible' to work, which is all that we require.

We now define \emph{heat} (sometimes written as $\dQ$) to be the one-form,
\begin{equation}\label{eq:heat}
  \xi := dU + \omega,
\end{equation}
Although the forms $\omega$, $\xi$, and $U$ are defined on $L$, they can equally be viewed as forms on the complete thermodynamic state space $M$, which can be pulled back to forms on $L$ via the embedding $\varphi:L\to M$.

As in ordinary gauge theory, a \emph{gauge symmetry} on the work line bundle is a smooth bundle automorphism $\Phi\in\Aut_N(L)$, characterised by a shift of each point $p\in L$ to another point on the same fibre, which thus satisfies,
\begin{equation}
  \pi(\Phi(L))=\pi(L).
\end{equation}
Gauge symmetries are fibre-wise affine functions that transform the vertical energy function $U$ as,
\begin{align}
  U\mapsto a U + b,
\end{align}
where $a=a_N\circ\pi$ and $b=b_N\circ\pi$ for some smooth functions $a_N:N\to\RR$ and $b_N:N\to\RR$. Thus, thermodynamic gauge symmetry is just the formal rescaling of units of energy, but which preserves the separation of energy into operationally accessible and inaccessible parts.

\section{Heat as a gauge connection}

A smooth subbundle $H \subseteq TL$ is called an \emph{connection} (or \emph{horizontal distribution}) for the fibre space $\pi:L\to N$ if,
\begin{equation}\label{eq:}
  T_pL = H_p \oplus V_p
\end{equation}
for all $p\in L$, where $V = \ker\pi_*$ is the space of vertical vectors. Thus, $H$ assigns to each $p$ a linear subspace $H_p$ with $H_p\cap V_p = \{0\}$, and such that the pushforward $\pi_*$ maps $H_p$ isomorphically onto $T_{\pi(p)}N$. This notion of a connection is sometimes called an \emph{Ehresmann connection} \cite{hermann1975g}. It is a generalisation of the standard notion of an affine connection on a principal bundle.

The distribution $H=\ker\xi$ is physically relevant for thermal processes because it defines what it means to be thermally insulated or `adiabatic'. Namely, a curve $\gamma:[0,1]\to L$ is called an \emph{adiabat} if there is no heat along its tangent vector field, $\bar{\gamma}\in\ker\xi$, where $\bar{\gamma}$ is the tangent vector field of $\gamma$. We will refer to $H$ as the \emph{adiabatic connection}. This terminology is justified by:

\begin{lemm}\label{lemm:1}
  Let $\pi:L\rightarrow N$ be a line bundle, $U$ a vertical coordinate function, and $\omega$ a one-form on $L$ such that $\ker\pi_*\subseteq\ker\omega$ and $\xi := dU + \omega$ is nowhere-vanishing. Then $H=\ker\xi$ is a connection.
\end{lemm}
\begin{proof}
Let $\partial_U$ be a vertical vector field satisfying $dU(\partial_U)=1$. Then $\partial_U\in \ker\pi_* \subseteq \ker\omega$, and so,
\begin{equation}
  \xi(\partial_U) = (dU + \omega)(\partial_U) = dU(\partial_U) = 1.
\end{equation}
Thus, the only element of $\ker\pi_*$ that is also in $\ker\xi$ is the zero vector, $\ker\pi_*\cap\ker\xi = \{0\}$. Moreover, since $\xi$ is nowhere-vanishing, we have $\dim\ker\xi=n-1$, and since $\pi$ is a line bundle, $\dim\ker\pi_*=1$. Combining these facts we get a decomposition $T_pL = H_p \oplus \left.\ker\pi_*\right|_p$ at every point $p\in L$.
\end{proof}
The key ingredient needed to treat heat as a connection is thus dimension-counting: the kernel of a nowhere-vanishing one-form $\xi$ is always $(n-1)$-dimensional. Thus, when the vertical fibres are one-dimensional and transverse to $\xi$, they together provide a basis for the $n$-dimensional tangent space at each point.

\section{Equilibrium as zero curvature}

In gauge physics, the curvature of the connection encodes the local forces and interactions on a bundle, with vanishing curvature corresponding to their absence. The situation is similar in thermodynamics, but with vanishing curvature corresponding to a system in `local equilibrium' in the sense of locally admitting an entropy and temperature function. This follows from two further lemmas, each of which is well known but which are rarely combined.

First, a smooth distribution $H\subseteq TL$ is called \emph{integrable} if every point $p\in L$ is contained in a non-empty immersed \emph{integral submanifold} $\varphi:S\to L$, defined to be a submanifold satisfying $T_qS=H_q$ for all $q\in S$. Given a fibre bundle $\pi:L\to N$ with a connection $H$ and $TL = H\oplus V$ for some vertical subbundle $V$, the \emph{curvature} of $H$ is the map $\Omega:H\times H\to V$ given by,
\begin{equation}\label{eq:curvature}
  \Omega(X,Y) := \mathrm{vert}[X,Y],
\end{equation}
where `$\mathrm{vert}$' denotes the projection onto $V$. This definition of curvature follows Hermann \cite{hermann1975g}, and is essentially the same as the standard definition in terms of a Lie algebra-valued one-form on a principal $G$-bundle (cf. Corollary 5.3 of \cite{KobayashiNomizu1963fdg}).

Curvature vanishes for all $X,Y\in H$ if and only if $[X,Y]\in H$ (Theorem 5.2 of \cite{KobayashiNomizu1963fdg} and Theorem 2.1 of \cite{hermann1975g}). In other words, $\Omega$ vanishes everywhere if and only if $H$ is \emph{involutive}. By the celebrated Frobenius theorem, involutivity is equivalent to integrability, and so we have:
\begin{lemm}\label{lemm:2}
  A connection on a fibre bundle has vanishing curvature if and only if it is integrable.
\end{lemm}
\begin{proof}
  By definition, $\Omega(X,Y)=0$ for all $X,Y\in H$ if and only if the $H$ is involutive as a distribution, which by the Frobenius theorem is equivalent to integrability.
\end{proof}

Next, in thermodynamics, the integrability of heat is equivalent to the existence of entropy and temperature functions \cite{boyling1968c}, which we will record here as:
\begin{lemm}\label{lemm:3}
  Let $\xi$ be a nowhere-vanishing one-form on a manifold $L$. Then $\ker\xi$ is integrable if and only if $\xi=TdS$ for some smooth $T,S$ in a neighbourhood of every point.
\end{lemm}
\begin{proof}
Integrability is equivalent to the existence of a local foliation by integral submanifolds, with a local defining function $S$ whose level sets are the leaves of the foliation. Thus, $\ker\xi$ is integrable only if there is a local function $S$ such that $\ker\xi = \ker dS$, which means that $\xi$ and $dS$ are proportional with respect to a smooth non-vanishing function $T$.
\end{proof}

Combining all three of our lemmas we now have:
\begin{thm}\label{thm}
Let $\pi:L\rightarrow N$ be a line bundle, $U$ a vertical coordinate function, and $\omega$ a one-form on $L$ with $\ker\pi_*\subseteq\ker\omega$ and $\xi := dU + \omega$ nowhere-vanishing. Then $H=\ker\xi$ is a connection, and the following are equivalent:
  \begin{enumerate}
    \item[$(A)$] $\xi = TdS$ for some smooth functions $T,S$, in a neighbourhood of every point.
    \item[$(B)$] The curvature of $H$ vanishes everywhere.
  \end{enumerate}
\end{thm}
\begin{proof}
  That $H$ is a connection is established by Lemma 1, while the equivalence of $(A)$ and $(B)$ is established by Lemmas 2 and 3.
\end{proof}
Thus, viewing adiabats as the lines of parallel transport of a gauge connection defined by heat, we find that the connection has vanishing curvature exactly when entropy and temperature are locally defined.

\section{Holonomy structure}

The fact that entropy and temperature arise when the adiabatic connection has vanishing curvature has further physical significance: in this section we will show that it is equivalent to having a trivial holonomy group, and in the next that it is equivalent to a conservation law.

Given a fibre bundle $\pi:L\to N$ and a point $p\in L$, a connection $H$ associates each curve $\gamma_N:[0,1]\to N$ beginning at $\pi(p)$ with a \emph{horizontal lift}, defined to be the unique curve $\gamma:[0,1]\to L$ such that $\pi(\gamma(t))=\gamma_N(t)$ for all $t\in[0,1]$ and $\bar{\gamma}\subset H$. When $\gamma_N$ is a closed loop, $q=\gamma_N(0)=\gamma_N(1)$, the horizontal lift defines a smooth map from the vertical fibre over $q$ to itself, called a \emph{holonomy}. The group of all such maps under composition is called the \emph{holonomy group} of $H$ over $q$. There is a well developed theory of the relationship between holonomy and curvature of connections on principal $G$-bundles, such as the famous Ambrose-Singer theorem. However, an important case for thermodynamics is the trivial case of vanishing curvature, which can be established directly in a more general context:

\begin{lemm}\label{lemm:B-C}
Let $\pi:L\rightarrow N$ be a fibre bundle and $H$ a connection. Then the following are equivalent:
  \begin{enumerate}
    \item[$(B)$] The curvature $\Omega$ of $H$ vanishes everywhere.
    \item[$(C)$] $H$ defines a trivial holonomy group in some neighbourhood of every point.
  \end{enumerate}  
\end{lemm}
\begin{proof}
Curvature in our sense is a tensor, and is thus determined entirely by its value at each point, unlike the Lie bracket. Let $X_p,Y_p \in H_p$ for some $p\in L$. Let $\tilde{X}_q := \pi_*X_p$ and $\tilde{Y}_q := \pi_*Y_p$ with $q:=\pi(p)$. By choosing appropriate local coordinates, we can always extend $\tilde{X}_q, \tilde{Y}_q$ to coordinate vector fields $\tilde{X}$ and $\tilde{Y}$ satisfying $[\tilde{X},\tilde{Y}]=0$ in a neighbourhood of $q$. Their horizontal lifts $X,Y$ are then vector fields extending $X_p,Y_p$ in a neighbourhood of $p$, and with the property that $[X,Y]$ is purely vertical, since $\pi_*[X,Y]=[\pi_*X,\pi_*Y]=0$ and the restriction of $\pi_*$ to $H$ is a linear isomorphism onto $TN$. Thus,
\begin{equation}
  \Omega(X,Y)|_p=\mathrm{vert}[X,Y]|_p=[X,Y]|_p.
\end{equation}
We will now show that $[X,Y]_p=0$ if and only if the holonomy group is locally trivial.

For any $t\geq 0$, let $\phi^{\tilde{X}}_t$ and $\phi^{\tilde{Y}}_t$ be the flows generated by $\tilde{X}$ and $\tilde{Y}$, respectively, and consider the sequence of flows,
\begin{equation}
  \tilde{\sigma}_{t} := \phi^{\tilde{Y}}_{-t}\circ \phi^{\tilde{X}}_{-t}\circ \phi^{\tilde{Y}}_{t}\circ \phi^{\tilde{X}}_{t}.
\end{equation}
Since $[\tilde{X},\tilde{Y}]=0$, these flows commute. Thus, for each value of $t$, the  $\tilde{\sigma}_t(q)=q$ traces out a piecewise-smooth rectangle that is closed.

By our assumption, the holonomy group is assumed to be trivial in a neighbourhood of every point. So, when the value of $t$ is sufficiently small, the horizontal lift of the closed rectangle is itself closed. Write $\phi^{X}_t$ and $\phi^{Y}_t$ for the flows generated by $X$ and $Y$, the sequence of flows defining the lifted rectangle is $\sigma_t = \phi^{Y}_{-t}\circ \phi^{X}_{-t}\circ \phi^{Y}_{t}\circ \phi^{X}_{t}$. A well known result \citep[Theorem 4.12]{frankel2012gop} now establishes that, for every smooth function $f$,
\begin{equation}
  [X,Y](f)\big|_p = \lim_{t\to 0} \frac{f(\sigma_t(p)) - f(p)}{t^2}.
\end{equation}
The right-hand side vanishes, since for all $t$ we have $\sigma_t(p)=p$. But, $f$ was arbitrary, so $[X,Y]|_p = 0$. Thus, the vertical component vanishes, which implies that $\Omega(X,Y)|_p=0$.
\end{proof}

When the holonomy group is non-trivial, the horizontal lift of a closed curve is not necessarily closed, but may spiral up to a new point on the fibre. However, it follows immediately from our definitions that the spiral does close when the holonomy group is trivial: 

\begin{lemm}\label{lemm:C-D}
Let $\pi:L\rightarrow N$ be a fibre bundle and $H$ a connection. Then the following are equivalent:
  \begin{enumerate}
    \item[$(C)$] $H$ defines a trivial holonomy group in some neighbourhood of every point.
    \item[$(D)$] The adiabatic lift of every closed curve is closed in some neighbourhood of every point.
  \end{enumerate}
\end{lemm}
The implication of Lemmas \ref{lemm:B-C} and \ref{lemm:C-D} for our purposes may be summarised as follows.
\begin{coro}\label{coro}
  Under the conditions of Theorem 1, the following are equivalent:
\begin{enumerate}
  \item[$(A)$] $\xi = TdS$ for some smooth functions $T,S$, in a neighbourhood of every point.
  \item[$(B)$] The curvature of $H$ vanishes everywhere.
  \item[$(C)$] $H$ defines a trivial holonomy group in some neighbourhood of every point.
  \item[$(D)$] The adiabatic lift of every closed curve is closed in some neighbourhood of every point.    
\end{enumerate}
\end{coro}
The `neighbourhood' restriction here indicates that this perspective on entropy and temperature is local, just like Carath\'eodory's derivation of entropy and temperature \cite{bernstein1960c,boyling1968c}.

\section{Conservation and Jauch's conjecture}

Jauch \cite{baronjauch1972a,jauch1972th,jauch1975at} claimed that entropy and temperature can be derived from a hypothesis of conservation, in a fashion similar to the standard derivation of the first law. His conservation hypothesis can be summarised: \emph{if a process produces no heat and restores all the work configurations to their initial values, then the process produces no work either}. The idea is that, unless there is some further source of work energy that has not been accounted for, the total work during such a cycle must vanish.

Let $L$ be an $n$-manifold, with $U:L\to\RR$ denoting energy, $\omega=\sum_{i=1}^{n-1}P_idV_i$ denoting work, and $\xi=dU + \omega$ denote heat. Jauch's hypothesis is,
\begin{enumerate}
  \item[\emph{if} (a)] $\gamma:[0,1]\to L$ is a closed adiabat such that $V_i(\gamma(0)) = V_i(\gamma(1))$ for each $i=1,...,{n-1}$, \emph{then}
  \item[(b)] $\int_\gamma\omega = 0$.
\end{enumerate}
Jauch conjectured that when his hypothesis holds, heat can be locally written as $\xi = TdS$ \cite{jauch1972th}. Unfortunately, his argument appears to contain an error in his Equation (2), which assumes that every point $p$ admits a neighbourhood in which it is connected to every other point by an adiabat. That is not generally true, and indeed contradicts Carath\'eodory's principle.

Nevertheless, an alternative proof is available, which proceeds by translating Jauch's hypothesis into the language of gauge theory. As before, let us view energy, work, and heat as defined on the total space $L$ of a line bundle $\pi:L\to N$. Then,
\begin{enumerate}
  \item[(a')] $V_i(\gamma(0)) = V_i(\gamma(1))$ for each $i=1,\dots,n-1$ means that the projection $\pi[\gamma]$ is a closed curve in $N$. And, the assertion that $\gamma$ is an adiabat ($\bar{\gamma}\in\ker\xi$) says that it is the horizontal lift of that closed curve defined by $H=\ker\xi$.
  \item[(b')] $\int_\gamma\omega = 0$ is equivalent to the statement that $\gamma$ is closed, since $\int_\gamma\omega = \int_\gamma(\xi - dU) = -\int_\gamma dU$ when $\xi(\bar{\gamma})=0$, which vanishes if and only if $U(\gamma(0))=U(\gamma(1))$.
\end{enumerate}
Therefore, Jauch's hypothesis is equivalent to our claim $(D)$, that the horizontal lift of every closed curve on $N$ is closed, and so a proof of his conjecture is a special case of our Corollary \ref{coro}, namely $(D)\Rightarrow(A)$.

\section{Thermal geometric phase}

Viewing heat as a gauge connection highlights two distinct ways in which a system can be out of equilibrium. First, it can fail to be in thermal equilibrium locally, which occurs when thermal curvature fails to vanish, or equivalently when Jauch's conservation hypothesis fails. Second, it can fail to be in thermal equilibrium globally, even when it is in equilibrium locally. This can happen through a thermal analogue of geometric phase \cite{ShapereWilczek1989}. 

Thermal geometric phase can be probed experimentally: consider a system $\pi:L\to N$ whose work configuration space is a circle, $N=S^1$. For example, $N$ could describe the angle $\theta$ of the key in a wind-up clock. Or, it could be the angle of a rotary piston like the Wankel engine that is thermally isolated, has no combustion, and has no intake or exhaust. The work done by this system on its environment is,
\begin{equation}
  \omega = -\tau d\theta,
\end{equation}
where $\tau:L\to\RR$ is a pressure function. Let $(U,\theta)$ provide a coordinate system for $L$. Locally, meaning for small changes in $\theta$, we will assume that $\tau$ depends only on $\theta$. Then, when rotating the system through a small angle $\theta$, define
\begin{equation}
  S := U - \int_{0}^\theta \tau(\theta)d\theta.
\end{equation}
If we further set $T := 1$, then $dU + \omega = \xi = TdS$, and so the connection $H=\ker\xi$ is flat and in local equilibrium. Nevertheless, a complete revolution through $\theta$ certainly does not restore energy to its initial value. Instead, total energy increases: the wind-up clock will develop tension, and the Wankel engine will develop pressure. Viewing heat as a gauge connection, this change in energy through a complete revolution is identical to a change in geometric phase.

\section{Conclusion}

We have seen that when thermodynamics is viewed as a gauge theory, heat defines a gauge connection associated with adiabatic processes. The existence of local entropy and temperature functions such that heat is equal to $TdS$ occurs exactly in the `interaction-free' case of vanishing curvature. However, this does not necessarily give rise to global equilibrium, due to the possible presence of a thermal analogue of geometric phase. This suggests a natural converse perspective: certain gauge theories display structures that are the same as those required for thermodynamics. However, whether such similarities are more than an analogy may depend on further physical considerations.

\subsection*{Acknowledgements} This research was funded by the European Research Council (ERC) Horizon Grant 101171014.

\end{document}